2.Two pendulum of length 1.21m and 1m start vibrating. At some instant the two are in the mean position in same phase. After how many vibrations of the longer pendulum will they be in the same phase again?
\

Grade:12th pass

2 Answers

Abhishek Singh
9 years ago
since time period depends on sq. root of length
T1/T2=sqr. root(L1/L2)
        = 1.1
it imples T1=1.1T2   ------>     10T1=11T2
it implies after 10 s\oscillations of first pendulum and 10 of the second , they will be again in phase.
 
Yash Chourasiya
3 years ago
Dear Student

IfT1​be the time period of pendulum 1(l=1.21m)andT2​that of the second one(l=1m).
We can say T1/T2 =√L1/L2
= 1.1 = 11/10
i.e. 10T1​ = 11T2​
That means10oscillations ofT1​is completed in the same time in which11oscillations ofT2​is completed. Hence after that time they again be in the same phase, that is, swinging together.

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya